a rectangular piece of cardboard is 4times as long as it is wide. you make a box by cutting squares with side length x from the corners of the cardboard and turning up the resulting flaps. write an expression involving x and the original width of the cardboard w that represents the perimeter of the base of the box. then simplify the expression. explain what the possible values of x can be.
Answer by asim(1)

width=w
length=4w
side of square=x
after turning the sides
the length of base=4w-x
width=w-2x
perimeter= 2(L+W)
= 2(4w-2x+w-2x)
= 2(5w-4x)
=10w-8x
possible values of x
10w-8x>0
10w>8x
10w/8>x
5w/4>x
or
x<5w/4